OpenStudy (anonymous):

how do you do - (5z + 12) =18

OpenStudy (anonymous):

first move the minus sign to the other side

OpenStudy (anonymous):

you get 5z+12=-18 then move the 12 to the other side 5z=-18-12 therefore 5z=-30 therefore z=-6

OpenStudy (anonymous):

hope this helps sweetheart!